# Speaker Anonymisation for Speech-based Suicide Risk Detection

### v1

@@ Abstract
Adolescent suicide is a critical global health issue, and speech provides a cost-effective modality for automatic suicide risk detection. Given the vulnerable population, protecting speaker identity is particularly important, as speech itself can reveal personally identifiable information if the data is leaked or maliciously exploited. This work presents the first systematic study of speaker anonymisation for speech-based suicide risk detection. A broad range of anonymisation methods are investigated, including techniques based on traditional signal processing, neural voice conversion, and speech synthesis. A comprehensive evaluation framework is built to assess the trade-off between protecting speaker identity and preserving information essential for suicide risk detection. Results show that combining anonymisation methods that retain complementary information yeilds detection performance comparable to that of original speech, while achieving protection of speaker identity for vulnerable populations.
@@ 1 Introduction
Suicide is a critical global health challenge and is one of the leading causes of death among adolescents [1, 2]. Early detection of suicide risk is essential for effective prevention and intervention of potential suicide attempts. Existing approaches still face limitations: clinical interviews require well-trained professionals, while self-report questionnaires are vulnerable to bias, discrimination, and even intentional concealment [3]. Speech serves as a cost-effective, remote, and non-invasive candidate for automatic suicide risk detection [4]. Speech naturally encodes both semantic and paralinguistic information. As an individual becomes pre-suicide, their speech exhibits measurable alterations in acoustic properties (e.g., jitter, shimmer, fundamental frequency, formant structure) [5], increased disfluency (e.g., hesitations and speech errors) [6], and distinct lexical patterns [7]. Previous studies have shown that suicide risk is detectable from spontaneous speech [8, 4].
Despite its great potential, it is critical to emphasise the privacy protection of participants’ speech data, especially given the dual vulnerabilities of age and psychological distress in the relevant population. As illustrated in Fig. 1, directly incorporating the original speech into the automatic detection system could pose serious privacy risks, as the speaker’s identity can be easily revealed if the data is leaked or maliciously exploited. By contrast, applying proper anonymisation before exposing speech to the system makes it much harder to trace the processed audio back to individuals, even in the event of a data breach. In fact, privacy protection has long been a central concern in the healthcare research, where anonymisation has been extensively studied for textual records [9], medical imaging [10], and physiological signals such as ECG [11] and EEG [12], where techniques range from simple de-identification [9, 12] to advanced generative models [10, 11]. Speech data poses unique challenges: it simultaneously encodes linguistic, acoustic, and emotional information, many of which are relevant for clinical analysis but also carry identifiable speaker characteristics [13], making it essential to carefully evaluate anonymisation techniques on both effectiveness in suppressing speaker identity and the ability to preserve features relevant to downstream suicide risk detection.
@@ 2.1 Speech-based Suicide Risk Detection
We perform speech-based suicide risk detection using original and anonymised speech respectively to examine whether using anonymised speech leads to significant performance degradation.
@@ 2.3 Anonymisation Performance Evaluation Framework
We build a multi-dimensional evaluation framework to comprehensively assess the effectiveness of different anonymisation methods, which spans over the following five critical aspects: speech quality, speaker discrimination, deviation in fundamental frequency contour, preservation of semantic content, and preservation of emotional content.
@@
Preservation of Semantic Content. The semantic consistency is evaluated by character error rate (CER) between the original and anonymised speech, both transcribed by the Paraformer ASR system.
@@ 4.2 Suicide Risk Detection Using Anonymised Speech
Fig. 4 shows the suicide risk detection results using speech anonymised by different approaches. The original speech got an average accuracy of 0.702, serving as the uppe-bound. Among all anonymisation methods, RVC produced the strongest single-model performance with an accuracy of 0.680, only 2% drop compared to original audio, while maintaining high anonymisation strength (with EER of 0.510). CosyVoice achieved the second-best result (accuracy of 0.658, 5% decay), possibly benefiting from the best preservation of semantic content while suffering from the loss of acoustic and prosodic cues that are important for suicide risk detection.
We further ensemble the CosyVoice and RVC models by averaging their predicted probabilities. This ensemble benefits from the complementary strengths of the two approaches where CosyVoice retains most semantic contents and RVC maintains most F0 contour. The combined system achieves an accuracy of 0.692. This result surpasses all individual systems, reduces the gap to the original speech to only 1%, while still ensuring good anonymisation. These findings suggest that combining complementary anonymisation strategies enhances speech utility while preserving privacy.

### v2

@@ Abstract
Adolescent suicide is a critical global health issue, and speech provides a cost-effective modality for automatic suicide risk detection. Given the vulnerable population, protecting speaker identity is particularly important, as speech itself can reveal personally identifiable information if the data is leaked or maliciously exploited. This work presents the first systematic study of speaker anonymisation for speech-based suicide risk detection. A broad range of anonymisation methods are investigated, including techniques based on traditional signal processing, neural voice conversion, and speech synthesis. A comprehensive evaluation framework is built to assess the trade-off between protecting speaker identity and preserving information essential for suicide risk detection. Results show that combining anonymisation methods that retain complementary information yields detection performance comparable to that of original speech, while achieving protection of speaker identity for vulnerable populations.
@@ 1 Introduction
Suicide is a critical global health challenge and one of the leading causes of death among adolescents [16, 19]. Early detection of suicide risk is essential for effective prevention and intervention of potential suicide attempts. Existing approaches still face limitations: clinical interviews require well-trained professionals, while self-report questionnaires are vulnerable to bias, discrimination, and even intentional concealment [5]. Speech serves as a cost-effective, remote, and non-invasive candidate for automatic suicide risk detection [24]. Speech naturally encodes both semantic and paralinguistic information. As an individual becomes pre-suicide, their speech exhibits measurable alterations in acoustic properties (e.g., jitter, shimmer, fundamental frequency) [3], increased disfluency (e.g., hesitations and speech errors) [21], and distinct lexical patterns [1]. Previous studies have shown that suicide risk is detectable from spontaneous speech [24, 2].
Despite its great potential, it is critical to emphasise the privacy protection of participants’ speech data, especially given the dual vulnerabilities of age and psychological distress in the relevant population.
As illustrated in Fig. 1, directly incorporating the original speech into the automatic detection system could pose serious privacy risks, as the speaker’s identity can be easily revealed if the data is leaked or maliciously exploited. By contrast, applying proper anonymisation before exposing speech to the system makes it much harder to trace the processed audio back to individuals, even in the event of a data breach. In fact, privacy protection has long been a central concern in the healthcare research, where anonymisation has been extensively studied for textual records [10], medical imaging [8], and physiological signals such as ECG [15] and EEG [13], where techniques range from simple de-identification [10, 13] to advanced generative models [8, 15]. Speech data poses unique challenges: it simultaneously encodes linguistic, acoustic, and emotional information, many of which are relevant for clinical analysis but also carry identifiable speaker characteristics [22], making it essential to carefully evaluate anonymisation techniques on both effectiveness in suppressing speaker identity and the ability to preserve features relevant to downstream suicide risk detection.
@@ 2.1 Speech-based Suicide Risk Detection
We perform speech-based suicide risk detection using original and anonymised speech respectively to examine whether using anonymised speech leads to remarkable performance degradation. A two-sided t-test between the results obtained with original and anonymised speech is conducted to assess statistical significance.
@@ 2.3 Anonymisation Performance Evaluation Framework
We build a multi-dimensional evaluation framework to comprehensively assess the effectiveness of different anonymisation methods, spanning over the following five critical aspects: speech quality, speaker discrimination, deviation in fundamental frequency, preservation of semantic content, and preservation of emotional content.
@@
Preservation of Semantic Content. The semantic consistency is evaluated by character error rate (CER) between the original and anonymised speech, both transcribed by Paraformer ASR system.
@@ 4.2 Suicide Risk Detection Using Anonymised Speech
Fig. 4 shows the suicide risk detection results using speech anonymised by different approaches. The original speech got an average accuracy of 0.702, serving as the upper-bound. Among all anonymisation methods, RVC produced the strongest single-model performance with an accuracy of 0.680, only 2% drop compared to original audio (with p=0.349), while maintaining high anonymisation strength (with EER of 0.510). CosyVoice achieved the second-best result (accuracy of 0.658, 5% decay, with p=0.136), possibly benefiting from the best preservation of semantic content while suffering from the loss of acoustic and prosodic cues. All other approaches show significant performance decay.
We further ensemble the CosyVoice and RVC models by averaging their predicted probabilities. This ensemble benefits from the complementary strengths of the two approaches where CosyVoice retains most semantic contents and RVC maintains most F0 contour. The combined system achieves an accuracy of 0.692. This result surpasses all individual systems, reduces the gap to the original speech to only 1% (with p=0.677), while still ensuring good anonymisation. These findings suggest that combining complementary anonymisation strategies enhances speech utility while preserving privacy.